OpenStudy (anonymous):

The volume of a cylinder is 81 cubic feet. What is the volume of a cone that fits exactly inside the cylinder? Provide an explanation and proof for your answer to receive full credit.

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (cwrw238):

vol of cylinder = pi r^2 h , vol of cone = (1/3) pi r^2 h so volume of the cone = 1/3 * 81 = 27 cu ft

OpenStudy (anonymous):

Well, the volume of a cone is 1/3 the volume of a cylinder. So what is 1/3 of 81?

OpenStudy (anonymous):

its 27

OpenStudy (anonymous):

i just need help walking through it

OpenStudy (anonymous):

i dont understand how to answer it

OpenStudy (midhun.madhu1987):